Write an equation of the line containing the given point and parallel to the given line.
(3,-2); 8x-7y=4
Thank you!

Answer by ikleyn(52780)

.
The left side of your equation will be EXACTLY THE SAME as of the given equation:


8x - 7y = c.       (*)


You need to determine the constant "c" in the right side.


For it, use the condition that your line contains the given point.


In other words, substitute the coordinates of the given point into equation (*) to find "c":


8*3 - 7*(-2) = 24 + 14 = 38.


Thus c = 38 and the equation the problem is asking for is 


8x - 7y = 38.

Solved.